# NeurDB: On the Design and Implementation of an AI-powered Autonomous Database

### v1

@@ Abstract.
Databases are increasingly embracing AI to provide autonomous system optimization and intelligent in-database analytics, aiming to relieve end-user burdens across various industry sectors. Nonetheless, most existing approaches fail to account for the dynamic nature of databases, which renders them ineffective for real-world applications characterized by evolving data and workloads. This paper introduces NeurDB, an AI-powered autonomous database that deepens the fusion of AI and databases with adaptability to data and workload drift. NeurDB establishes a new in-database AI ecosystem that seamlessly integrates AI workflows within the database. This integration enables efficient and effective in-database AI analytics and fast-adaptive learned system components. Empirical evaluations demonstrate that NeurDB substantially outperforms existing solutions in managing AI analytics tasks, with the proposed learned components more effectively handling environmental dynamism than state-of-the-art approaches.
@@ 1. Introduction
Database management systems (DBMSs) are becoming more automated and intelligent by embracing artificial intelligence (AI). AI empowers these systems to achieve autonomous optimization Kraska et al. (2019); Pavlo et al. (2017); Zhu et al. (2024) and support complex analytics Karanasos et al. (2020); Xing et al. (2024); Ooi et al. (2015) with minimal human intervention, thereby catering to advanced data- and AI-centric applications.
@@
The aspiration of integrating DBMSs with AI was first expressed forty years ago Brodie (1988), which has been periodically revisited with evolving technology Wang et al. (2016); Cong et al. (2024); Ré et al. (2015). With the advancements in both AI and DBMSs, considerable progress has been made in deepening their fusion. Unfortunately, a huge gap remains between the potential of this integration and its current state of usability, largely due to the inherent differences in their paradigms. The dynamic nature of databases, characterized by constantly changing data and workloads due to transactional updates, and user interactions, poses a fundamental challenge. For example, an e-commerce database might experience a sharp increase in workloads during flash sales compared to normal periods, caused by huge volumes of updates from sales transactions and processing. In contrast, AI models typically derive intelligence from static datasets and thus can become outdated quickly in the face of database dynamism. For example, models for learned query optimization often acquire knowledge by training on specific system environments including workload patterns and data distributions. However, when these environments drift, e.g., due to data updates from user interactions, models relying on past knowledge may lead to ineffective query optimizations since they are not updated to reflect new conditions.
The discrepancy in dynamicity between AI and DBMSs poses a great challenge to the effective adoption of AI-enhanced DBMSs. First, traditional AI-driven system optimization, including knob tuning Pavlo et al. (2017), intelligent advisors Zhu et al. (2024); Li et al. (2021), and learned system components Kraska et al. (2019) such as query optimizers Marcus et al. (2021); Zhu et al. (2023), indexes Kraska et al. (2018), and concurrency control Wang et al. (2021), often overlook the dynamicity of data distributions and workloads. Second, another important facet of AI and DBMS integration, namely in-database AI analytics for complex tasks such as disease progression predictions and user purchase recommendations, has gained traction but faces similar adaptability issues. Existing systems generally focus on static model inference Salazar-Díaz et al. (2024); Park et al. (2022), model serving Vartak (2017), and others Xu et al. (2022), without considering necessary model adaptations to evolving data and workload patterns. Consequently, they require frequent manual updates and retraining to remain effective, which complicates the analytics workflow. Recent works Wu and Ives (2024); Li et al. (2022, 2023a); Kurmanji and Triantafillou (2023) aim to address these challenges but are typically more effective in controlled scenarios, such as data distribution changes following certain rules or query optimization under read-only workloads. These methods have not been designed to generalize across more dynamic and practical workloads involving continuous read and write operations.
@@
We envision a deep integration of AI and DBMSs for adaptability to both data and workload drift. Unlike existing approaches that merely overlay AI onto DBMSs or selectively enhance certain system components, we aim for a comprehensive fusion of AI with DBMSs that enables continuous adaptation to drift. However, achieving this level of integration presents significant challenges. First, the integration necessitates a fundamental redesign of the entire AI workflow, including training, inference, and fine-tuning, within the database architecture. Second, models for learned database components and AI analytics must swiftly adapt to drifting data workloads without losing effectiveness and efficiency. Therefore, developing such a system requires both a system foundation that seamlessly supports model adaptation and AI models that are inherently adaptive in structure.
@@
In this paper, we present NeurDB [^2] https://github.com/neurdb/neurdb, an AI-powered autonomous DBMS that provides efficient and effective in-database AI analytics to seamlessly support modern AI-powered applications, and fully embraces AI techniques in each major system component to offer self-driving capabilities. At the core of NeurDB is an in-database AI ecosystem that achieves deep integration of the AI workflow into the database. In this ecosystem, we develop multiple in-database AI operations, such as model training, inference, and fine-tuning, along with an in-database AI engine to handle the execution of these operators. NeurDB can then directly support in-database AI analytics by calling these AI operations. To further simplify end-users in submitting their complex AI analytics tasks to NeurDB, we provide a user-friendly interface by extending SQL syntax with PREDICT syntax. Moreover, we propose two techniques to optimize the AI ecosystem performance. First, we devise a data streaming protocol that significantly reduces data transfer overhead, yielding better performance for AI operations. Second, we develop an incremental update technique to minimize the fine-tuning cost, facilitating the fast adaptation of AI models. Based on this foundation, we propose two efficient learned system components, namely a learned concurrency control algorithm and a learned query optimizer, that can adapt quickly to data and workload drift.
2. Preliminaries
@@
In this section, we outline the design goals of NeurDB and present the SQL syntax enhanced for AI-powered Analytics.
@@ 2.1. Design Goals
DBMSs are dynamic as data distributions and workloads evolve over time, and therefore, the system must be designed for adaptability while also guaranteeing reliability and scalability. Building upon this understanding, we define three essential properties for NeurDB.
@@
Informally, the design goal of NeurDB is to ensure the system can uphold reliability as quickly as possible when the system or its components adapt or evolve due to data and workload drift, while ensuring good scalability.
@@ 2.2. SQL Syntax for AI Analytics
NeurDB incorporates enhanced SQL, which extends from the standard SQL to support AI analytics. As illustrated in Listing 1 and Listing 2, NeurDB introduces a new PREDICT keyword to handle two typical ML tasks: classification with the CLASS OF clause and regression with the VALUE OF clause. Inspired by the original principle of introducing SQL that allows an application developer to write a query with SELECT, and then let the DBMS find the most efficient way to store and retrieve data, we ensure that a developer can submit an AI analytic task simply with PREDICT. All the following operations, such as retrieving training data and invoking AI models, are handled automatically by the database. We also plan to expose more enhanced SQL for AI model management and other AI analytics functionalities. Due to space constraints, we mainly focus on the PREDICT query in this paper.
We now present two real-world analytic scenarios that can be directly supported by NeurDB using PREDICT.
@@
The disease progression prediction in healthcare can be handed by the SQL shown in Listing 1. By specifying the features with TRAIN ON and using data stored in the ‘diabetes’ table and directly input with VALUES, we predict the class of the ‘outcome’ variable, which indicates whether a patient has diabetes or not.
Regression
Listing 2 shows the SQL to predict and fill in the missing scores of products based on user reviews. Similarly, it specifies the features for training and the data table to predict the value of the ‘score’ variable. Notably, it automatically excludes features with unique constraints to avoid including meaningless data.
Listing 1: Syntax for a classification task
@@
Listing 2: Syntax for a regression task
1PREDICT VALUE OF score
2FROM review
3WHERE brand_name = ’SpecialGoods’
4TRAIN ON * WITH brand_name <> ’SpecialGoods’
@@ 3. System Overview
We now describe the system architecture of NeurDB as shown in Figure 1, where we achieve the deep fusion of AI and databases by establishing an in-database AI ecosystem. To achieve this, we first holistically redesign each existing database component, such as the query optimizer, transaction manager, and index manager, to support in-database AI analytics and function as learned components for improved performance. In particular, we enhance the SQL parser and the query optimizer to produce a customized query plan for PREDICT queries. Beyond traditional operators for fetching and processing data, e.g., scan and join, the query executor in NeurDB includes in-database AI operators for model training, inference, and fine-tuning. With these AI operators, we effectively align the AI workflow with the database query processing. We are developing additional AI operators, such as model selection, to provide more comprehensive AI services. For example, a query may call the model selection operator (denoted as MSelection) to automatically select the best-suited model for a given prediction task Xing et al. (2024), thereby enhancing accuracy and efficiency. We then introduce a new component, called the AI engine, into NeurDB. It handles the AI-related processing requests from AI operators and learned components and creates AI tasks on their behalf, where the task manager dispatches them to the CPU/GPU runtime for execution. Specifically, we propose a data streaming protocol to reduce data transfer overhead, and an incremental model update technique to facilitate fast model adaptation, thus enhancing the performance of the AI engine. We shall elaborate on these optimizations in Section 4.1. Also, we design dedicated AI model storage to store AI models and serve them based on requests from the AI engine. We further implement a monitor to detect unexpected performance or accuracy issues, based on which we trigger automatic and appropriate model adaptation.
In Figure 1, we illustrate a running example of a PREDICT query. After parsing and optimizing the query, the executor performs the scan operation to retrieve data and then invokes the inference operator to deliver the inference task to the AI engine. Next, the AI engine uploads the model from the model storage if it is not in the model buffer, and conducts model inference to produce the results.
@@ 4. The Design of NeurDB
In this section, we present two key subsystems of NeurDB, namely an in-database AI ecosystem and learned system components.
@@ 4.1. In-database AI Ecosystem
Model Manager
Given the dynamic nature of DBMSs, a typical AI lifecycle extends beyond a single model. As new data is introduced, incremental updates and retraining are required to address data and workload drifts that degrade predictive performance. This process results in multiple evolving model versions, creating significant management difficulties and storage overheads. To address this challenge, NeurDB leverages the capabilities of databases to efficiently manage AI models and handle drift by design. Specifically, it introduces a dedicated model manager in its AI engine, enabling fine-grained model management with efficient updates. The model manager provides high-level interfaces for handling AI operations, such as training, inference, and fine-tuning, executed via model views. Similar to data views in DBMSs, model views serve as logical abstractions of AI models tailored for specific tasks, with physical representations maintained in model storage. Formally, a model $M$ is uniquely defined by its Model ID (MID) $i$ and the timestamp of creation $t$, i.e., $M_{i,t}$. A model $M_{i,t}$, created by either users or internal components, comprises a series of layers $L$, each identified by a Layer ID (LID) $j$ and its training timestamp. To generate outputs for data $X$, the layers are executed sequentially: $M_{i,t}(X)=L^{(n)}(L^{(n-1)}(...(L^{(1)}(X)))$. This layered model storage approach aligns with the structure of deep neural networks (DNNs), ensuring fast and efficient AI model accessibility. Notably, this model formulation can also be generalized to non-linear DNN architectures that are represented as directed acyclic graphs, which can be achieved by executing layers based on the topological order.
@@
Leveraging its layered model storage, NeurDB enables linear versioning for a specific model and supports incremental model updates through fine-tuning. In particular, to adapt the current model to drifting data distributions, the AI engine selectively fine-tunes the final layers using the updated data stored in the designated database relation and freezes the preceding layers. Subsequently, only the updated layers remain persistent in the model storage, from where they can be extracted and merged with the previously frozen layers to create a new model version.
Figure 3 illustrates how the model manager supports incremental updates for AI models. Whenever an inference or fine-tuning task is initiated, the system retrieves all layers up to the most recent version to construct the model, i.e., for a model $M_{i,t}$ with $k$ layers, $M_{i,t}(X)=L_{i,t_{k}}^{(k)}(L_{i,t_{k-1}}^{(k-1)}(...(L_{i,t_{1}}^{(1)}(X))))\;s.t.\;\forall p,q\in\{1,...,k\},L_{i,t_{p}}^{(p)}\rightarrow t_{p}\geq t_{q}\wedge t_{p}\leq t.$ Let us consider the example shown in Figure 3. Supposing the model $M_{1}$ requires fine-tuning the last layer $L_{n}$, $M_{1,2}$, namely the second version of $M_{1}$, can then be assembled by layers $\{L_{1,1}^{(1)},...,L_{1,1}^{(n-1)},L_{1,2}^{(n)}\}$. This allows $M_{1,1}$ and $M_{1,2}$ to share the majority of model weights, ensuring adaptability to data drift while maintaining efficient storage.

### v2

@@ Abstract.
Databases are increasingly embracing AI to provide intelligent in-database analytics and autonomous system optimization, aiming to relieve end-user burdens across various industry sectors. Nonetheless, most existing approaches fail to account for the dynamic nature of databases, which renders them ineffective for real-world applications characterized by evolving data and workloads. This paper introduces NeurDB, an AI-powered autonomous database that deepens the fusion of AI and databases with adaptability to data and workload drift. NeurDB establishes a new in-database AI ecosystem that seamlessly integrates AI workflows within the database. This integration enables efficient and effective in-database AI analytics and fast-adaptive learned database components. Empirical evaluations demonstrate that NeurDB substantially outperforms existing solutions in managing AI analytics tasks, with the proposed learned components more effectively handling data and workload drift than state-of-the-art approaches.
@@ 1. Introduction
Database management systems (DBMSs) are becoming more intelligent and autonomous by embracing artificial intelligence (AI), catering to modern data- and AI-centric applications. AI augments DBMS functionality to support in-database AI analytics Karanasos et al. (2020); Xing et al. (2024); Ooi et al. (2015); Wang et al. (2015), enabling complex and advanced analytics tasks such as disease progression predictions and user purchase recommendations. In parallel, AI empowers DBMSs to achieve autonomous DBMS optimization, driving a broad range of learned database designs that enhance system performance with minimal human intervention, including learned knob tuners Pavlo et al. (2017), learned system drivers Zhu et al. (2024); Li et al. (2021), and learned database components Kraska et al. (2019) such as query optimizers Marcus et al. (2021); Zhu et al. (2023), indexes Kraska et al. (2018), and concurrency control Wang et al. (2021), etc.
@@
The aspiration of integrating DBMSs with AI was first expressed forty years ago Brodie (1988), which has been periodically revisited with evolving technology Wang et al. (2016); Cong et al. (2024); Ré et al. (2015). With the advancements in both AI and DBMSs, considerable progress has been made in deepening their fusion. Unfortunately, a huge gap remains between the potential of this integration and its current state of usability, largely due to the inherent differences in their paradigms Ooi et al. (2024). The dynamic nature of databases, characterized by data and workload drift, poses a fundamental challenge Wu and Ives (2024); Kurmanji et al. (2024); Kurmanji and Triantafillou (2023). For example, suppose an e-commerce database stores user and product information. While the database continuously evolves due to user and merchant activities in normal periods, it would experience a sharp increase in workloads during flash sales, where huge volumes of updates from sales transactions are processed. In contrast, AI models typically derive intelligence from static datasets and thus can become outdated quickly in the face of database dynamism. For instance, models for user purchase recommendations typically acquire knowledge by training on specific user behaviors and product information. However, when the data drifts due to transactional updates, models relying on past knowledge may produce inaccurate recommendations since they are not updated to reflect new conditions. Similarly, learned query optimizers trained based on historical system environments including data distributions and workload patterns can struggle to deliver effective query plans as continuous data and workload drift.
Adapting to data and workload drift has gained traction for both in-database AI analytics and autonomous DBMS optimization. Existing databases supporting in-database AI analytics SQL (2024); PostgresML (2024); MindsDB (2024) often recommend users to completely retrain models with new data and workloads when the models become outdated. This retraining process is typically performed manually outside the database, and therefore, it may complicate the AI analytics workflow and becomes inefficient when dealing with continuous data and workload drift. To tackle this problem, in the context of autonomous DBMS optimization, recent works Wu and Ives (2024); Li et al. (2022, 2023a); Kurmanji and Triantafillou (2023) have started addressing the adaptability issues by automatically detecting drift and triggering model retraining. However, these approaches are often tailored to certain components. For example, approaches Wu and Ives (2024); Kurmanji and Triantafillou (2023) that allow learned query optimizers to withstand workload pattern changes can be ineffective for learned indexes requiring rapid adaptation to data distribution drift. Consequently, a paradigm that robustly handles data and workload drift for various learned database components remains absent.
@@
We envision a deeper integration of AI and DBMSs for enhanced adaptability to both data and workload drift. Unlike existing approaches that merely overlay AI onto DBMSs or selectively enhance certain system components, we aim for a comprehensive fusion of AI with DBMSs that enables continuous adaptation to drift. It promises a cohesive AI-powered DBMS that provides advanced in-database AI analytics and deeply integrates AI into all its key components. However, achieving this level of integration presents significant challenges. First, the integration necessitates a fundamental redesign of the entire AI workflow, including training, inference, and fine-tuning, within the database architecture. Second, models for AI analytics and learned database components must swiftly adapt to data and workload drift without losing effectiveness and efficiency. Therefore, developing such an AI-powered database requires both a system foundation that seamlessly supports model adaptation and AI models that are inherently adaptive in structure.
@@
In this paper, we present NeurDB, an AI-powered autonomous DBMS that provides efficient and effective in-database AI analytics to seamlessly support modern AI-powered applications, and fully embraces AI techniques in each major system component to offer autonomous system optimization. At the core of NeurDB is an in-database AI ecosystem that deeply integrate the AI workflow into the database. In this ecosystem, we develop multiple in-database AI operators, such as model training, inference, and fine-tuning, along with an in-database AI engine to handle the execution of these operators. NeurDB can then directly support in-database AI analytics by calling these AI operators. To further simplify end-users in submitting their AI analytics tasks to NeurDB, we provide a user-friendly interface by extending SQL with PREDICT syntax. Moreover, we propose two techniques to optimize the AI ecosystem performance. First, we devise a data streaming protocol that significantly reduces data transfer overhead, yielding better performance for AI operations. Second, we develop an incremental update technique to minimize the fine-tuning cost, facilitating the fast adaptation of AI models. Based on the in-database AI ecosystem, we enable two efficient learned database components, namely a learned concurrency control algorithm and a learned query optimizer, that can adapt quickly to data and workload drift.
The remainder of the paper is structured as follows. The next section provides the necessary preliminaries and further clarifies the motivation. Section 3 describes the system overview of NeurDB, and Section 4 details the key system design, including the in-database AI ecosystem and fast-adaptive learned database components. Section 5 presents the experimental results. Section 6 discusses the related works before Section 7 concludes.
2. Preliminaries and Motivation
@@
In this section, we introduce the preliminaries of AI and DBMS integration, outline the design goals of NeurDB, and present the enhanced SQL syntax designed for in-database AI analytics.
2.1. AI and DBMS Integration
The integration of AI and DBMSs can offer each other mutual benefits Wang et al. (2016), and building on this intuition, various research works on AI and DBMS integration have been proposed, which can be broadly categorized into two areas: in-database analytics and autonomous system optimization. Modern databases, such as Oracle Oracle (2024), Microsoft Azure SQL SQL (2024), Amazon Redshift Redshift (2024), and Google BigQuery BigQuery (2024), support in-database AI analytics, and allow AI tasks to be performed directly through user-friendly SQL interfaces. For better performance, these databases enable AI-driven knob tuning to autonomously refine database configurations.
Adapting to data and workload drift has been recognized as a critical challenge in advancing the integration of AI and DBMSs Wu and Ives (2024); Kurmanji et al. (2024); Kurmanji and Triantafillou (2023). Many existing approaches for whether in-database analytics Karanasos et al. (2020); Salazar-Díaz et al. (2024) or autonomous system optimization Marcus et al. (2019); Kraska et al. (2018) depend on static models or complete retraining, limiting their ability to handle continuous data and workload drift. More importantly, these approaches are typically layered on top of DBMSs rather than being holistically embedded within DBMSs. This separation hinders real-time adaptability, because internal and fine-grained performance metrics, which are essential for detecting and addressing drift, can be difficult to capture externally Wang et al. (2021); Li et al. (2022).
To mitigate these limitations, we design and develop NeurDB, an AI-powered database with AI holistically embedded from the ground up. By deeply integrating AI into all its key components and functionalities, NeurDB aims to automatically manage adaptation workflows for both in-database AI analytics and autonomous system optimization.
@@ 2.2. Design Goals
DBMSs are dynamic as data and workloads evolve over time, and therefore, the system must be designed for adaptability while also guaranteeing reliability and scalability. Building upon this understanding, we equip NeurDB with these three key properties.
@@
Informally, our design goal is to ensure NeurDB can uphold reliability as quickly as possible when the system adapts or evolves due to data and workload drift, while ensuring good scalability. In addition, we design NeurDB as a general-purpose AI-powered DBMS which is expected to serve as a foundational infrastructure for advanced data- and AI-centric applications to enrich various domains with AI capabilities Ooi et al. (2024).
@@ 2.3. SQL Syntax for AI Analytics
NeurDB incorporates enhanced SQL to support AI analytics. As illustrated in Listing 1 and Listing 2, it extends from the standard SQL by introducing a PREDICT keyword to handle two typical AI tasks: regression with the VALUE OF clause and classification with the CLASS OF clause. Inspired by the original principle of SQL that allows an application developer to write a query with SELECT, and then let the DBMS find the most efficient way to store and retrieve data, we ensure that a developer can submit an AI analytics task simply with PREDICT. All the following operations, such as retrieving training data and invoking AI models, are handled automatically by NeurDB. Consequently, unlike existing solutions PostgresML (2024); BigQuery (2024) that require application developers to intervene in the execution of AI analytics, such as specifying model parameters, our approach can make the execution transparent to application developers, relieving them from the complexities of performing such analytics. Further, we plan to expose additional SQL for AI model management and other AI analytics functionalities, such as generative AI.
We now present two real-world analytics scenarios that can be directly supported by NeurDB using PREDICT queries.
Regression
Listing 1 shows the SQL to predict and fill in the missing scores of products based on user reviews. By specifying the features for training with TRAIN ON and using data stored in the ‘review’ table, we predict the value of the ‘score’ variable. Notably, for TRAIN ON, the asterisk automatically excludes features with unique constraints to avoid including meaningless data.
Listing 1: Syntax for a regression task
1PREDICT VALUE OF score
2FROM review
3WHERE brand_name = ’SpecialGoods’
4TRAIN ON *
5 WITH brand_name <> ’SpecialGoods’
@@
The disease progression prediction in healthcare can be handed by the SQL shown in Listing 2. Similarly, this query specifies the features for training with TRAIN ON and the data table using FROM to predict the class of the ‘outcome’ variable, which indicates whether a patient has diabetes or not. In addition, we support directly inputting missing data using VALUES.
Listing 2: Syntax for a classification task
@@ 3. System Overview
We now describe the system architecture of NeurDB as shown in Figure 1, where we achieve the deep fusion of AI and databases by establishing an in-database AI ecosystem. Based on this AI ecosystem, we enable efficient and effective in-database AI analytics and develop fast-adaptive learned database components.
In-database AI Ecosystem
We establish the in-database AI ecosystem by holistically redesigning existing database components, such as the query executor, and introducing additional modules, including an AI engine and AI model storage. We first extend the query executor to support AI workflow. Beyond traditional operators for fetching and processing data, e.g., scan and join, the query executor in NeurDB includes in-database AI operators for model training, inference, and fine-tuning. With these AI operators, we effectively integrate the AI workflow into the database query processing. We are developing additional AI operators, such as model selection Xing et al. (2024) and model slicing Zeng et al. (2025), to provide more comprehensive AI services. For example, a query may call the model selection operator (denoted as MSelection) to automatically select the best-suited model for a given prediction task, thereby enhancing accuracy and efficiency. We then introduce a new module, called the AI engine, into NeurDB. It handles the AI-related processing requests from AI operators and learned components and creates AI tasks on their behalf, where the task manager dispatches them to the CPU/GPU runtime for execution. Specifically, we propose a data streaming protocol to reduce data transfer overhead, and an incremental model update technique to facilitate fast model adaptation, thus enhancing the performance of the AI engine. We shall elaborate on these optimizations in Section 4.1. Also, we design dedicated AI model storage to store AI models and serve them based on requests from the AI engine. We further implement a monitor to detect unexpected performance or accuracy issues, based on which we trigger automatic and appropriate model adaptation.
In-database AI Analytics
In Figure 1, we illustrate a running example of a PREDICT query. We enhance the SQL parser and the query optimizer to produce a customized query plan for PREDICT queries. After parsing and optimizing the query, the query executor then executes the query according to the query plan. In particular, it performs the scan operator to retrieve data and then invokes the inference operator to deliver the inference task to the AI engine. Next, the AI engine uploads the model from the model storage if it is not in the model buffer, and conducts model inference to produce the results. If the model is detected to be inaccurate, NeurDB invokes the fine-tuning operator to update the outdated model with the help of the AI engine.
Autonomous DBMS Optimization
We also enable autonomous system optimization based on the proposed in-database AI ecosystem. For instance, the learned query optimizer interacts with the AI engine to train a model responsible for generating efficient query plans. The monitor tracks the performance of the generated query plans, and if a plan is identified as inefficient, the monitor notifies the AI engine to fine-tune the model, enabling continuous optimization and adaptability. Similarly, the monitor can trigger autonomous knob tuning when suboptimal knob settings are detected, ensuring that the system remains well-configured to handle data and workload drift effectively. We will further introduce the detailed design of learned database components in Section 4.2.
@@ 4. The Design of NeurDB
In this section, we present two key subsystems of NeurDB, namely an in-database AI ecosystem and learned database components.
@@ 4.1. In-database AI Ecosystem
Model Manager
Given the dynamic nature of DBMSs, a typical AI lifecycle extends beyond a single model. As new data is introduced, incremental updates and retraining are required to address data and workload drift that degrade predictive performance. This process results in multiple evolving model versions, creating significant management difficulties and storage overheads. To address this challenge, NeurDB leverages the capabilities of databases to efficiently manage AI models created by either users or internal components and handle drift by design. Specifically, it introduces a dedicated model manager in its AI engine, enabling fine-grained model management with efficient updates. The model manager provides high-level interfaces for handling AI operations, such as training, inference, and fine-tuning, executed via model views. Similar to data views in DBMSs, model views serve as logical abstractions of AI models tailored for specific tasks, with physical representations maintained in model storage. Formally, a model $M$ comprises a series of layers $L^{(j)}(j\in\mathbb{Z}^{+})$. To generate outputs for data $X$, the layers are executed sequentially, i.e., $M(X)=L^{(n)}(L^{(n-1)}(...(L^{(1)}(X)))$. This layered model storage approach aligns with the structure of deep neural networks (DNNs), ensuring fast and efficient AI model accessibility. Notably, this model formulation can also be generalized to non-linear DNN architectures that are represented as directed acyclic graphs, which can be achieved by executing layers based on the topological order.
@@
Leveraging its layered model storage, NeurDB enables versioning for a specific model and supports incremental model updates through fine-tuning. In particular, to adapt the current model to drifting data distributions, the AI engine selectively fine-tunes the final layers using the updated data stored in the designated database relation and freezes the preceding layers. Subsequently, only the updated layers remain persistent in the model storage, from where they can be extracted and merged with the previously frozen layers to create a new model version.
Figure 3 illustrates how the model manager supports incremental updates for AI models. Whenever an inference or fine-tuning task is initiated, the system retrieves all layers up to the most recent version to construct the model, i.e., for a model $M_{i,t}$, uniquely defined by its Model ID (MID) $i$ and the timestamp of creation $t$, consisting of $k$ layers,
$M_{i,t}(X)=L_{i,t_{k}}^{(k)}(L_{i,t_{k-1}}^{(k-1)}(...(L_{i,t_{1}}^{(1)}(X))))$
$\;s.t.\;\forall p,q\in\{1,...,k\},L_{i,t_{p}}^{(p)}\rightarrow t_{p}\geq t_{q}\wedge t_{p}\leq t.$
Let us consider the example shown in Figure 3. Supposing the model $M_{1}$ requires fine-tuning the last layer $L_{n}$, $M_{1,2}$, namely the second version of $M_{1}$, can then be assembled by layers $\{L_{1,1}^{(1)},...,L_{1,1}^{(n-1)},L_{1,2}^{(n)}\}$. This allows $M_{1,1}$ and $M_{1,2}$ to share the majority of model weights, ensuring adaptability to data drift while maintaining efficient storage.